OpenStudy (anonymous):

There are (42)3 ⋅ 40 horses on a stud farm. What is the total number of horses on the farm? 0 46 47 424

OpenStudy (anonymous):

@Michele_Laino

Nnesha (nnesha):

first of all use `^` for exponent example 2 to the 3rd power 2^3

OpenStudy (anonymous):

yeah i know that

Nnesha (nnesha):

\[\huge\rm (4^2)^3 \times 4^0\] so it should looks like this

Nnesha (nnesha):

two exponent rules to solve that question 1st) anything to the zero power equal to one \[\huge\rm (x)^0 = 1 ~~~~(Anything)^0 =1\] and 2nd )\[\huge\rm (a^m)^n=a^{m \times n}\]

Nnesha (nnesha):

so 4^0 = ??

Nnesha (nnesha):

how would you simplify (4^2)^3 = ?

OpenStudy (michele_laino):

hint: using the rule of power of power and the rule of multiplication of power wityh the same basis, we can write: \[\Large {\left( {{4^2}} \right)^3} \cdot {4^0} = {4^{2 \cdot 3 + 0}} = ...?\]

OpenStudy (michele_laino):

with*

OpenStudy (anonymous):

4^6

OpenStudy (michele_laino):

correct!